# Force-free wave interaction in magnetar magnetospheres: Computational modeling in axisymmetry

### v1

@@ Acknowledgements
The authors thank Andrei Beloborodov, Alex Chen, Amir Levinson, Alexander Philippov, Juan de Dios Rodríguez, Lorenzo Sironi, Anatoly Spitkovsky, Yajie Yuan, and Muni Zhou for useful discussions. We appreciate the support of the Department of Energy (DoE) Early Career Award DE-SC0023015 (PI: L. Sironi). This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470, PI: L. Sironi). This research was facilitated by the Multimessenger Plasma Physics Center (MPPC), NSF grant PHY-2206609. JFM acknowledges support from the National Science Foundation (NSF) under grant No. AST-1909458 (PI: A. Philippov) and AST-1814708 (PI: A. Spitkovsky). MAA acknowledges support from grant PID2021-127495NB-I00, funded by MCIN/AEI/10.13039/501100011033 and by the European Union under ‘NextGenerationEU’ as well as ‘ESF: Investing in your future’. Additionally, MAA acknowledges support from the Astrophysics and High Energy Physics program of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union ‘NextGenerationEU’ (PRTR-C17.I1) as well as support from the Prometeo excellence program grant CIPROM/2022/13 funded by the Generalitat Valenciana. JFM thanks the Department of Astronomy at Tsinghua University (Beijing, China) for their hospitality during a collaborative visit stimulating this research. The presented simulations were enabled by the Ginsburg cluster (Columbia University Information Technology), the MareNostrum supercomputer (Red Española de Supercomputación, AECT-2022-3-0010), and the Frontera supercomputer (Stanzione et al., 2020) at the Texas Advanced Computing Center (LRAC-AST21006). Frontera is made possible by the NSF award OAC-1818253.
@@ Appendix A Fit of the decaying FMS tail
Figure 10: Approximation of the decaying tail of FMS pulses for the wave-train discussed in Section 3.2 by a fit function (Equation A1). We indicate several fit parameters in the figure, including the decay onset time $x_{0}$ and the decay rate $\tau$ (following the dashed black line).
In Section 3.2 we discuss the ongoing production of low-frequency FMS waves by AWs bouncing in the inner magnetospheres. At later times, AWs are strongly affected by grid diffusion due to their extreme shear accumulating during the propagation along the diverging dipole field lines. Figure 4 summarizes the extreme requirements on the numerical resolution for accurately capturing the decay of the outgoing FMS wave pulses. Here, we discuss how the expected luminosity function can be parametrized with an appropriate fit to approximate the total energy stored in the decaying FMS tail. We propose to analyze the expression:
@@
Figure 10 shows a fit of Equation (A1) to the high-resolution ($N_{R_{\ast}}=1024$) data generated in Section 3.2. Our numerical optimization yields the following fit parameters and standard errors: $a_{0}=(6.61\pm 0.19)\times 10^{-4}$, $a_{1}=(0.38\pm 0.03)$, $x_{0}=(8.43\pm 0.01)$, $\tau=(2.80\pm 0.07)$, and $T_{b}=(3.71\pm 0.03)$. We can use these results to estimate the total energy $\tilde{E}_{\rm tail}/E_{\rm inj}\approx 0.16\%$ (Equation A2). Evaluating the fit function indicates that the modulation period $T_{b}$ of the decaying luminosity flux tail roughly corresponds to the AW crossing time of the dipolar field line of the initial injection colatitude, where $cT_{\rm bounce}/R_{\ast}\approx 3.5$. We finally note that the amount of energy transferred to FMS waves during the first peak and the first half of the second peak is comparable to the energy in the tail, so that the total energy transferred from AWs to FMS waves is $\Delta E_{\rm FMS}/E_{\rm inj}\sim 0.3\%$.
@@ Appendix B Scales of the outgoing FMS waves
Figure 11: Analysis of the scales of FMS waves generated by the conditions outlined in Section 3.2 (see Figure LABEL:fig:a1bounce). The left panel outlines the FMS wave along the equator with its high-frequency (red color) and low-frequency (blue color) components. We indicate the characteristic field line length for the propagation of seed AWs as a gray-shaded region. The right panel shows the power spectrum of the high-frequency FMS wave component (in arbitrary units). The average wave field is subtracted before the frequency analysis. The outgoing FMS waves have half the wavelength of the seed AWs.
This section analyzes the properties of the FMS waves induced by the bouncing train of AWs discussed in Section 3.2. For $ct/R_{\ast}=8.76$ (Figure LABEL:fig:a1bounce, third column), we assess the scales of the low-frequency and high-frequency wave components in Figure 11. The high frequency signal has a wavelength corresponding to half of the seed AW wavelength, with a clear peak in the Fourier analysis shown in the right panel of Figure 11. The variations of the low-frequency part correspond to the characteristic length $L\approx 3.5R_{\ast}$ of the field lines along which the seed AWs propagate.

### v2

@@ Acknowledgements
We thank the referee for feedback to our manuscript and the careful review of Section 2.2. The authors thank Andrei Beloborodov, Alex Chen, Amir Levinson, Alexander Philippov, Juan de Dios Rodríguez, Lorenzo Sironi, Anatoly Spitkovsky, Yajie Yuan, and Muni Zhou for useful discussions. We appreciate the support of the Department of Energy (DoE) Early Career Award DE-SC0023015 (PI: L. Sironi). This work was supported by a grant from the Simons Foundation (MP-SCMPS-00001470, PI: L. Sironi). This research was facilitated by the Multimessenger Plasma Physics Center (MPPC), NSF grant PHY-2206609. JFM acknowledges support from the National Science Foundation (NSF) under grant No. AST-1909458 (PI: A. Philippov) and AST-1814708 (PI: A. Spitkovsky). MAA acknowledges support from grant PID2021-127495NB-I00, funded by MCIN/AEI/10.13039/501100011033 and by the European Union under ‘NextGenerationEU’ as well as ‘ESF: Investing in your future’. Additionally, MAA acknowledges support from the Astrophysics and High Energy Physics program of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union ‘NextGenerationEU’ (PRTR-C17.I1) as well as support from the Prometeo excellence program grant CIPROM/2022/13 funded by the Generalitat Valenciana. JFM thanks the Department of Astronomy at Tsinghua University (Beijing, China) for their hospitality during a collaborative visit stimulating this research. The presented simulations were enabled by the Ginsburg cluster (Columbia University Information Technology), the MareNostrum supercomputer (Red Española de Supercomputación, AECT-2022-3-0010), and the Frontera supercomputer (Stanzione et al., 2020) at the Texas Advanced Computing Center (LRAC-AST21006). Frontera is made possible by the NSF award OAC-1818253.
@@ Appendix A Fit of the decaying FMS tail
Figure 10: Approximation of the decaying tail of FMS pulses for the wave train discussed in Section 3.2 by a fit function (Equation A1). We indicate several fit parameters in the figure, including the decay onset time $x_{0}$ and the decay rate $\tau$ (following the dashed black line).
In Section 3.2 we discuss the ongoing production of low-frequency FMS waves by Alfvén waves bouncing in the inner magnetospheres. At later times, Alfvén waves are strongly affected by grid diffusion due to their extreme shear accumulating during the propagation along the diverging dipole field lines. Figure 4 summarizes the extreme requirements on the numerical resolution for accurately capturing the decay of the outgoing FMS wave pulses. Here, we discuss how the expected luminosity function can be parametrized with an appropriate fit to approximate the total energy stored in the decaying FMS tail. We propose to analyze the expression:
@@
Figure 10 shows a fit of Equation (A1) to the high-resolution ($N_{R_{\ast}}=1024$) data generated in Section 3.2. Our numerical optimization yields the following fit parameters and standard errors: $a_{0}=(6.61\pm 0.19)\times 10^{-4}$, $a_{1}=(0.38\pm 0.03)$, $x_{0}=(8.43\pm 0.01)$, $\tau=(2.80\pm 0.07)$, and $T_{b}=(3.71\pm 0.03)$. We can use these results to estimate the total energy $\tilde{E}_{\rm tail}/E_{\rm inj}\approx 0.16\%$ (Equation A2). Evaluating the fit function indicates that the modulation period $T_{b}$ of the decaying luminosity flux tail roughly corresponds to the Alfvén wave crossing time of the dipolar field line of the initial injection colatitude, where $cT_{\rm bounce}/R_{\ast}\approx 3.5$. We finally note that the amount of energy transferred to FMS waves during the first peak and the first half of the second peak is comparable to the energy in the tail, so that the total energy transferred from Alfvén waves to FMS waves is $\Delta E_{\rm FMS}/E_{\rm inj}\sim 0.3\%$.
@@ Appendix B Scales of the outgoing FMS waves
Figure 11: Analysis of the scales of FMS waves generated by the conditions outlined in Section 3.2 (see Figure LABEL:fig:a1bounce). The left panel outlines the FMS wave along the equator with its high-frequency (red color) and low-frequency (blue color) components. We indicate the characteristic field line length for the propagation of seed Alfvén waves as a gray-shaded region. The right panel shows the power spectrum of the high-frequency FMS wave component (in arbitrary units). The average wave field is subtracted before the frequency analysis. The outgoing FMS waves have half the wavelength of the seed Alfvén waves.
This section analyzes the properties of the FMS waves induced by the bouncing train of Alfvén waves discussed in Section 3.2. For $ct/R_{\ast}=8.76$ (Figure LABEL:fig:a1bounce, third column), we assess the scales of the low-frequency and high-frequency wave components in Figure 11. The high-frequency signal has a wavelength corresponding to half of the seed Alfvén wave wavelength, with a clear peak in the Fourier analysis shown in the right panel of Figure 11. The variations of the low-frequency part correspond to the characteristic length $L\approx 3.5R_{\ast}$ of the field lines along which the seed Alfvén waves propagate.